A particle travels along a straight line with a velocity of v= (20 − 0.05s^2) m/s, where s is in meters. Determine the acceleration of the particle at s = 15 m.
1
Expert's answer
2020-08-24T13:21:03-0400

Velocity v=(20-0.05s"^2" )m/s

at s=15m ,

velocity v=(20-0.05"\times 15^2)"

v=18.75m/s

as we know acceleration is the rate of change of velocity

a="\frac{dv}{dt}"

="\frac{d(20-0.05s^2)}{dt}"

= 0-0.05"\times 2s\times \frac{ds}{dt}"

=-0.05"\times 2 \times 15\times v" ( since velocity is rate of change of distance)

= "-1.5\times18.75"

= - 28.125m/s"^2"

Hence the acceleration at s=15m is - 28.125m/s"^2"